# Searching for axion dark matter with magnetic resonance force microscopy

### v1

@@ Abstract
While dark matter is known to interact with baryonic matter gravitationally, ultra-weak non-gravitational couplings are present in most theoretical models. Many of these couplings have effects that manifest as a real or effective A.C. magnetic field oscillating at $\omega_{\rm DM}$. For instance, the axion-photon coupling and the dark photon induce effective currents that generate real electromagnetic fields sikivie1983experimental; wilczek1987two; arza2022earth; chaudhuri2015radio. Furthermore, the axion-fermion couplings induce effective magnetic fields on fermion spins barbieri1989axion; alexander2018detecting. Due to the wide viable mass range for ULDM around 20 orders of magnitude, various technologies are required to fully search the allowed parameter space. This underscores a critical need for the development of ultra-precise A.C. magnetometers across a vast frequency range for dark matter detection. Several dark matter searches are based on magnetic resonance techniques in magnetized samples crescini2018operation; crescini2020axion; flower2019broadening; ikeda2022axion; jackson2020overview; garcon2019constraints; walter2025search; xu2024constraining. Various other magnetometry techniques have been used to search for dark matter bloch2020axion; bloch2023constraints; lee2023laboratory; wei2025dark; bloch2022new; abel2023search; gao2022axion; karanth2023first; jiang2021search; gavilan2024searching. In this manuscript, we propose a new mechanical, frequency-tunable search for dark matter using magnetic resonance force microscopy (MRFM).
@@
Figure 1: Diagram of the MRFM setup. The top illustrates the general scenario where a magnetized sample under the influence of a D.C. field $B_{0}\hat{z}$ and a field due to a nearby micromagnet attached to a mechanical resonator undergoes magnetic resonance electron spins in the presence of a dark matter induced magnetic field $B_{\rm DM}\cos(\omega_{\rm DM}t)$. All the spins within the resonance region put a force on the resonator. A pump field $B_{\rm p}\cos(\omega_{\rm p}t)$ parallel to the $B_{\rm DM}$ amplifies the dark matter signal and modulates it down to the micromagnet-loaded resonator’s resonance frequency $|\omega_{\rm DM}-\omega_{\rm p}|=\omega_{\rm m}$, inducing vibrations that are read out optically. The bottom shows the two mechanical systems considered in this work: membranes and cantilevers.
@@
We first recast various dark matter searches via various couplings into an A.C. magnetic field detection problem. Then, we introduce a magnetic resonance force microscopy sidles1991noninductive; sidles1995magnetic; suter2004magnetic detection scheme involving frequency down-conversion and optimize it for the axion-electron coupling. We show that the axion-electron coupling can be probed beyond current laboratory constraints with existing technology in under a minute of integration time. We then discuss the frequency scanning capabilities of our proposed setup. Finally, we explore this setup’s capability to put constraints on other dark matter - Standard Model couplings.
@@
Dark Matter Induced Magnetic Fields.– Here we describe various ULDM couplings to Standard Model fields that manifest as effective or real magnetic field effects. Starting with axions, which are pseudoscalar particles that can be described by a classical wave of amplitude $\sqrt{2\rho_{\rm DM}}/\omega_{\rm DM}$ oscillating at its Compton frequency $\omega_{\rm DM}$. The axion field $a(\vec{x},t)$ can interact with fermions $\psi_{f}$ via the operator $g_{af}\partial_{\mu}a\bar{\psi}_{f}\gamma^{\mu}\gamma_{5}\psi_{f}$, where $g_{af}$ is the axion-fermion coupling strength, $\gamma^{\mu}$ are the Dirac matrices and $f=e,p,n$ for electrons, protons and neutrons. This leads to an interaction Hamiltonian of the form $\gamma_{f}\vec{S}\cdot\vec{B}_{\rm eff}$, where $\gamma_{f}=g_{f}e/2m_{f}$ is the fermion gyromagnetic ratio with the Lande g gactor $g_{f}$, $\vec{S}=\frac{\hbar}{2}\vec{\sigma}$ with Pauli matrices $\vec{\sigma}$, and $\vec{B}_{\rm eff}$ is an induced effective magnetic field barbieri1989axion; alexander2018detecting. For axion couplings to electrons, protons and neutrons, respectively, the effective magnetic field strengths $B_{af}$ are given below in terms of the coupling strength $g_{af}$:
@@
Axions can also couple to photons and induce an effective current that generates real magnetic fields sikivie1983experimental; wilczek1987two; caputo2021dark; arza2022earth. Here, a background magnetic field $B_{0}$ is required to mediate the interaction, and the induced magnetic field as a function of coupling strength $g_{a\gamma}$ is
@@
where $L$ is the size of the experiment. It is noteworthy that the axion-induced magnetic fields given above are independent of dark matter mass/frequency.
In addition to axions, dark photons coupling to Standard Model photons can also induce an effective current that generates real electromagnetic fields caputo2021dark; arza2022earth. The magnetic field generated by the dark photon as a function of the kinetic mixing parameter $\epsilon$ and experiment size $L$ is
@@
Unlike the axion couplings, the induced magnetic field here is dependent on dark matter mass/frequency. As before, these magnetic field strengths are written to show the required magnetometer sensitivity to match current laboratory constraints. While better astrophysical constraints may exist for certain couplings, these may be more model dependent, necessitating new laboratory searches to supplement these constraints. It is also worth reiterating that these equations give strengths of dark matter-induced magnetic fields $B_{\rm DM}$ which oscillate at $\omega_{\rm DM}$. See Appendix A for further details of the various dark matter induced magnetic fields. We now show how these weak signals can be detected by re-purposing an existing technology.
Magnetic Resonance Force Microscopy.– Magnetic Resonance Force Microscopy is a scanned-probe technique that combines imaging capabilities of magnetic resonance imaging with resolution of atomic force microscopy hammel2007magnetic. Unlike conventional magnetic resonance detection, which inductively detects the electromagnetic signal of transverse magnetization, MRFM directly senses changes in longitudinal magnetization through a mechanical resonator sidles1993signal. MRFM experiments generally employ cyclical modulation of longitudinal magnetization $M_{z}$ at the mechanical resonance frequency, using techniques like amplitude modulation (AM) Suter2004; wago1997magnetic. An advantage of this mechanical readout is avoiding circuit noise and associated backaction damping. Appendix B provides a more detailed comparison of MRFM with other magnetic resonance imaging techniques.
A typical setup involves a micromagnet attached to the mechanical resonator placed directly above the sample, as shown in Fig. 1. For an extended sample that is placed at a large enough distance from the micromagnet, the micromagnet can be treated as a point dipole suter2002probe; sidles91, as detailed in Appendix B. The micromagnet produces a strong local magnetic field gradient that couples the resonator to the spins of the magnetized sample underneath it. As a result of this interaction, the resonator senses a force given by Suter2004,
@@
To induce magnetic spin resonance, a static magnetic field $B_{0}$ along the $z$ direction polarizes the sample, and an oscillating magnetic field $B_{\rm rf}$ is applied perpendicular to it. The bias field $B_{0}$ and the micromagnet’s magnetic field $B_{z}$ determine the Larmor frequency $\omega_{L}=\gamma(B_{0}+B_{z})$. Here, the role of the radio frequency field is played by the dark matter-induced magnetic field $B_{\rm DM}\cos(\omega_{\rm DM}t)$. To enhance the weak dark matter signal, a strong pump field $B_{\rm p}\cos(\omega_{\rm p}t)$ is applied parallel to $B_{\rm DM}$ with $B_{\rm p}\gg B_{\rm DM}$, in an approach similar to ruoso2016quax.
The magnetization density $\vec{m}(\vec{r},t)$ of the magnetized sample can be calculated using stationary solutions of the Bloch equations abragam1961principles,
@@
as long as $|\omega_{\rm DM}-\omega_{\rm p}|\ll\omega_{L}$ augustine2002transient; ruoso2016quax. Here, $m_{0}=n_{s}\mu_{B}\tanh(\mu_{B}B_{0}/k_{B}T)$ is the equilibrium magnetization density, with $n_{s}$ as the spin number density, $T_{1}$ and $T_{2}$ as the longitudinal and transverse relaxation times. Due to the nonlinearity of the Bloch equations, the superposition of the two radio frequency fields, $B_{\rm p}cos(\omega_{\rm p})+B_{\rm DM}cos(\omega_{\rm DM})$, produces an on-resonance ($\gamma B_{0}-\omega_{\rm DM}=0$) beat-note modulation of $m_{z}$,
@@
where $\omega_{\rm D}=|\omega_{\rm p}-\omega_{\rm DM}|$ being the detection frequency satisfying the condition $\omega_{\rm D}^{-1}<\min\left\{T_{1},T_{2}\right\}$. Expression 8 holds at resonance and for dark matter frequency around $\rm 1~GHz$ $(\omega_{L}=\omega_{\rm DM}=\rm GHz)$ where the dipole radiation damping effects, mentioned in ruoso2016quax, can be ignored. While a stronger pump field leads to larger amplification, one must ensure that the system is kept far from saturation $B_{\rm p}\ll 1/\gamma\sqrt{T_{1}T_{2}}$. For further details on the detection scheme, see Appendix B.
@@
The integral is performed over the resonant part of the sample volume, where the electron spins are in magnetic resonance as defined by the condition $\omega_{\rm DM}=\omega_{L}\pm\delta\omega$ with linewidth given by $\delta\omega=\frac{1}{T_{2}}$ far from saturation. Near the zero-probe-field resonance (ZPFR) regime, the entire sample can be within the resonance region suter2004magnetic, which gives a large force on the mechanical system, despite the weak field gradient. To determine the sensitivity of a MRFM setup to radio frequency field $B_{\rm DM}$, we now consider the contribution of various noise sources to the resonator motion.
Figure 2: Sensitivity plot of axion-electron coupling strength $g_{aee}$ as a function of axion mass $m_{\rm DM}$ (lower axis) and corresponding Compton frequency $f_{\rm DM}$ (top axis). A second y-axis gives the strength of the effective magnetic field induced on electrons by axions corresponding to the coupling strength $g_{aee}$. Our results for MRFM based axion search are given in red and blue for a membrane and cantilever, respectively. For each resonator, a 100 day long run at a single frequency is shown as well as a 100 day scan of minute long runs at neighboring frequencies, showing that one can choose to increase sensitivity or bandwidth. The QCD axion band is shown in yellow and existing constraints from astrophysical capozzi2020axion, space gondolo2009solar; aprile2022search and laboratory experiments terrano2015short; yan2019constraining are shown in gray. Also shown are narrowband constraints from other magnetic resonance based searches in green, labeled QUAX crescini2018operation; crescini2020axion, UWA flower2019broadening and QND ikeda2022axion. The top right panel shows the finer details of a single sensitivity line including the two sidebands at $\omega_{p}\pm\omega_{m}$. The bottom right panel shows the contributions from the back action noise (purple), thermal noise (yellow), and imprecision noise (green) to the total noise equivalent force (red).
Figure 3: Sensitivity plot of dark photon kinetic coupling parameter $\epsilon$ as a function of dark photon mass $m_{\rm DM}$ (lower axis) and corresponding Compton frequency $f_{\rm DM}$ (upper axis). Our results for MRFM based axion search are given in red and blue for a membrane and cantilever, respectively. For each resonator, a 100 day long run at a single frequency is shown as well as a 100 day scan of minute long runs at neighboring frequencies, showing that one can choose to increase sensitivity or bandwidth. The gray region represents constraints from astrophysical and cosmological observations assuming the dark photon makes up majority of dark matter, while dark green and light green are constraints from haloscope searches and rescaled limits from axion searches. See caputo2021dark for details of various constraints on the dark photon kinetic coupling.
@@
where the dark matter coherence time is $\tau_{c}\equiv\tau_{\rm DM}=\frac{c^{2}}{v_{\rm DM}^{2}f_{\rm DM}}=\frac{10^{6}}{f_{\rm DM}}$ for couplings proportional to the field and $\tau_{c}\equiv\tau_{\nabla a}=0.68\ \tau_{\rm DM}$ for couplings proportional to the gradient of the axion field $a$ barbieri2017searching. Here, NEF denotes the noise equivalent force, which corresponds to a signal to noise ratio of 1.
@@
DM Parameter Space.– Before providing specific details of the device used for our estimates, we now overview existing constraints on the ULDM parameter space to work out the required magnetic field sensitivity. The parameter space for axion-electron coupling with existing laboratory and astrophysical constraints is shown in Fig. 2. The most stringent constraint on the axion-electron coupling at $g_{ae}\sim 10^{-13}~\rm eV$ stems from the lack of observed reduction in the tip of the red giant branch (TRGB) luminosity in globular clusters, where axion bremsstrahlung from electrons would otherwise cause significant energy loss and cooling capozzi2020axion. Weaker constraints arise from solar axion searches, including limits on axion-induced solar energy losses from measurements of the neutrino flux gondolo2009solar and null detection of electron recoils from solar neutrinos aprile2022search. The best current laboratory constraint in the relevant mass range is due to torsion pendulum measurements of dipole-dipole forces between electrons terrano2015short at $7\times 10^{-9}~\rm eV$, while measurement of the anomalous magnetic moment of electrons yan2019constraining also gives weaker constraints.
Three major milestones in the search for axions via their electron coupling are probing the coupling better than torsion pendulum limits terrano2015short, observations of tip of the red giant branch (TRGB) luminosity in globular clusters capozzi2020axion, and probing the QCD axion bands for DFSZ and KSVZ-like axions. These correspond to magnetic field strengths of $3\times 10^{-16}~\rm T$, $5\times 10^{-21}~\rm T$, and $2\times 10^{-24}~\rm T$, respectively (see Fig. 2). The best current narrowband experimental constraints are from magnetic axion haloscopes that rely on detecting the axion-magnon coupling crescini2018operation; crescini2020axion; flower2019broadening; ikeda2022axion. A number of proposals relying on optimistic advancements in technology have been proposed using QND measurements of axion-magnon conversion chigusa2020detecting; mitridate2020detectability; berlin2024physical; catinari2025hunting; chang2025mosaic. Our MRFM proposal for axion searches is similar in technical requirements and detection sensitivity to the best existing laboratory constraints, including QUAX crescini2018operation; crescini2020axion. Unlike the aforementioned proposals, it does not require major technological advancements and the experiment can be performed with current technology.
@@
Using the parameters given in Table 1 and equation 11, we find a minimum magnetic field sensitivity at resonance of about $2.0\times 10^{-16}$ T ($1.0\times 10^{-17}$ T) for a minute (100 days) of integration time with a membrane and $2.4\times 10^{-16}$ T ($1.2\times 10^{-17}$ T) for a minute (100 days) of integration time with a cantilever. Using Equations 1 and 4, we obtain minimum values that can be probed by our setup of the axion-electron coupling and the dark photon kinetic mixing parameters, respectively. These are plotted in Figs. 2 and 3. For the axion-electron coupling, an MRFM experiment can search for axions beyond existing laboratory constraints. For the dark photon kinetic mixing, this system allows for filling in gaps between existing haloscope constraints.
@@
Radius $r_{\rm sample}$ $1~\rm mm$
Table 1: Experimental parameters used in our calculations of the sensitivity of MRFM to dark matter-induced magnetic fields. The values for the resonators are chosen in the range of parameters from scozzaro2016magnetic; Takahashi2018; zwickl2012progress for membrane and wago1998paramagnetic for cantilever.
The bottom right panel of Fig. 2 shows contributions to the noise equivalent force spectral density from back action noise (purple), thermal noise (yellow) and imprecision noise (green). The top right panel shows the corresponding plot for sensitivity to the axion-electron coupling, showing zoomed in part of both sidebands $\omega_{p}\pm\omega_{m}$. This two-sideband feature comes from the frequency up conversion scheme and has been seen by, for example, the AURIGA detector branca2017search. Our proposed experiment is limited by thermal noise over a relatively broad band with a linewidth of around $100~\rm Hz$ (see Fig. 5 in Appendix C).
@@
One of the limits of the setup is the size of the sample. There is a limit to how large a YIG sphere can be fabricated, and large spheres may be too sensitive to adverse effects, such as radiation damping. Since our proposed setup allows for a large sample area, which translates to increased sensitivity to lower magnetic field values (see Eq. 12), one can use multiple YIG spheres, as done by QUAX crescini2020axion. A potential problem here may be the transfer of photons from radiation damping from one sphere to its neighbor. However, this problem can be resolved by insulating the spheres. In our calculations, we take a configuration of 9 spheres (see Fig 4). The number of YIG spheres, i.e. the number of spins, is not a technical limitation of the proposal, but will be limited for any experiment by available resources.
The constraints presented here involve just a minute of integration time on existing technology, with the entire experimental setup shown in figure 1 taking $\sim$1 cm3 of volume. It is possible to fit multiple such setups inside a dilution refrigerator, improving the sensitivity by $\sqrt{N_{e}}$, for $N_{e}$ experimental setups, or scanning over a much larger bandwidth. A higher mechanical quality factor, as demonstrated on several platforms, can allow one to probe even weaker ULDM couplings. There is also potential for further multiplexing by using more magnetized samples, which only requires one to adjust the applied bias field to widen the resonance region. Furthermore, for axion-electron couplings in particular, flux concentrators can be used to amplify the signal gao2025focusing. Finally, our proposal involves standard quantum limit readout, a limit that has already been circumvented in several experimental platforms, e.g. xia2023entanglement; jia2024squeezing; mason2019continuous.
Conclusion and outlook. – While most MRFM research has pushed towards detecting single spin flips, our work motivates exploration of another regime, namely MRFM as an ultra-weak magnetometer. This requires using a frequency scanning technique that can be adapted to other optomechanical searches for dark matter manley2020searching; manley2021searching; hirschel2024superfluid. Our proposal for a magnetic resonance force microscope search for dark matter induced magnetic fields is just one thrust within the larger optomechanics for fundamental physics theme carney2021mechanical; moore2021searching; kalia2024ultralight; ahrens2025levitated; bose2025massive. Beyond dark matter, other interesting signals from gravitational waves, neutrinos, etc. can potentially be probed using optomechanical systems, paving the way for pushing forward technology and potentially discovering new physics.

### v2

@@ Abstract
While dark matter is known to interact with baryonic matter gravitationally, ultra-weak non-gravitational couplings are present in most theoretical models, some of which may manifest as a real or effective A.C. magnetic field oscillating at $\omega_{\rm DM}$. For instance, the axion-photon coupling and the dark photon induce effective currents that generate real electromagnetic fields [89, 101, 10, 33]. Furthermore, axion-fermion couplings induce effective magnetic fields on fermion spins [15, 5]. Due to the wide viable mass range for ULDM, various technologies are required to fully search the allowed parameter space. This underscores a critical need for the development of ultra-precise A.C. magnetometers across a vast frequency range for dark matter detection. Several dark matter searches are based on magnetic resonance techniques in magnetized samples [36, 35, 44, 56, 58, 47, 97, 103]. Various other magnetometry techniques have been used to search for dark matter [19, 21, 66, 98, 20, 2, 45, 63, 61, 48]. In this manuscript, we propose a new mechanical, frequency-tunable search for dark matter using magnetic resonance force microscopy (MRFM).
@@
Figure 1: Diagram of the MRFM setup. The top illustrates the general scenario where a magnetized electron spin sample under influence of D.C. field $B_{0}\hat{z}$ and field of a micromagnet attached to a mechanical resonator undergoes magnetic resonance in the presence of a dark matter induced field $B_{\rm DM}\cos(\omega_{\rm DM}t)$. All spins within the resonance region put a force on the resonator. A pump field $B_{\rm p}\cos(\omega_{\rm p}t)$ parallel to the $B_{\rm DM}$ amplifies the dark matter signal and modulates it down to the micromagnet-loaded resonator’s resonance frequency $|\omega_{\rm DM}-\omega_{\rm p}|=\omega_{\rm m}$, inducing vibrations enabling optical detection. The bottom shows the two resonators considered in this work: membranes and cantilevers.
@@
We first recast several ULDM searches into an A.C. magnetic field detection problem. Then, we introduce a magnetic resonance force microscopy [88, 86, 92] detection scheme involving frequency down-conversion and optimize it for the axion-electron coupling. We show that the axion-electron coupling can be probed beyond current laboratory constraints with existing technology in under a minute of integration time. We then discuss the frequency scanning capabilities of our proposed setup. Finally, we explore this setup’s capability to put constraints on other ULDM - Standard Model couplings.
@@
Dark Matter Induced Magnetic Fields.– Here we describe various ULDM couplings to Standard Model fields that manifest as effective or real magnetic field effects. Starting with axions, which are pseudoscalar particles that can be described by a classical wave of amplitude $\sqrt{2\rho_{\rm DM}}/\omega_{\rm DM}$ oscillating at its Compton frequency $\omega_{\rm DM}$. The axion field $a(\vec{x},t)$ can interact with fermions $\psi_{f}$ via the operator $g_{af}\partial_{\mu}a\bar{\psi}_{f}\gamma^{\mu}\gamma_{5}\psi_{f}$, where $g_{af}$ is the axion-fermion coupling strength, $\gamma^{\mu}$ are the Dirac matrices and $f=e,p,n$ for electrons, protons and neutrons. This leads to an interaction Hamiltonian of the form $\gamma_{f}\vec{S}\cdot\vec{B}_{\rm eff}$, where $\gamma_{f}=g_{f}e/2m_{f}$ is the fermion gyromagnetic ratio with the Lande g gactor $g_{f}$, $\vec{S}=\frac{\hbar}{2}\vec{\sigma}$ with Pauli matrices $\vec{\sigma}$, and $\vec{B}_{\rm eff}$ is an induced effective magnetic field [15, 5]. For axion couplings to electrons, protons and neutrons, respectively, the effective magnetic field strengths $B_{af}$ are given below in terms of the coupling strength $g_{af}$:
@@
Axions can also couple to photons and induce an effective current that generates real magnetic fields [89, 101, 29, 10]. Here, a background magnetic field $B_{0}$ is required to mediate the interaction, and the induced magnetic field as a function of coupling strength $g_{a\gamma}$ is
@@
where $L$ is the size of the experiment.
In addition to axions, dark photons coupling to Standard Model photons can also induce an effective current that generates real electromagnetic fields [29, 10]. The magnetic field generated by the dark photon as a function of the kinetic mixing parameter $\epsilon$ and experiment size $L$ is
@@
While better astrophysical constraints may exist for certain couplings, these may be more model dependent, necessitating new laboratory searches to supplement these constraints. See Appendix A for further details of the various dark matter induced magnetic fields. We now show how these weak signals can be detected by re-purposing an existing technology.
Magnetic Resonance Force Microscopy.– Magnetic Resonance Force Microscopy is a scanned-probe technique that combines imaging capabilities of magnetic resonance imaging with resolution of atomic force microscopy [52]. Unlike conventional magnetic resonance detection, which inductively detects the electromagnetic signal of transverse magnetization, MRFM directly senses changes in longitudinal magnetization through a mechanical resonator [87]. MRFM experiments generally employ cyclical modulation of longitudinal magnetization $M_{z}$ at the mechanical resonance frequency, using techniques like amplitude modulation (AM) [91, 96]. An advantage of this mechanical readout is avoiding circuit noise and associated backaction damping. Appendix B provides a more detailed comparison of MRFM with other magnetic resonance imaging techniques.
A typical setup involves a micromagnet attached to the mechanical resonator placed directly above the sample, as shown in Fig. 1. For an extended sample that is placed sufficiently far from the micromagnet, the micromagnet can be treated as a point dipole [90, 85], as detailed in Appendix B. The micromagnet produces a strong local magnetic field gradient that couples the resonator to magnetized spin sample underneath it. As a result, the resonator senses a force given by [91],
@@
To induce magnetic spin resonance, a static magnetic field $B_{0}$ along the $z$ direction polarizes the sample, and an oscillating magnetic field $B_{\rm rf}$ is applied perpendicular to it. The bias field $B_{0}$ and the micromagnet’s magnetic field $B_{z}$ determine the Larmor frequency $\omega_{L}=\gamma(B_{0}+B_{z})$. Here, the role of the radio frequency field is played by the dark matter-induced magnetic field $B_{\rm DM}\cos(\omega_{\rm DM}t)$. To enhance the weak dark matter signal, a strong pump field $B_{\rm p}\cos(\omega_{\rm p}t)$ is applied parallel to $B_{\rm DM}$ with $B_{\rm p}\gg B_{\rm DM}$, in an approach similar to [82].
The magnetization density $\vec{m}(\vec{r},t)$ of the magnetized sample can be calculated using stationary solutions of the Bloch equations [3],
@@
as long as $|\omega_{\rm DM}-\omega_{\rm p}|\ll\omega_{L}$ [12, 82]. Here, $m_{0}=n_{s}\mu_{B}\tanh(\mu_{B}B_{0}/k_{B}T)$ is the equilibrium magnetization density, with $n_{s}$ as the spin number density, $T_{1}$ and $T_{2}$ as the longitudinal and transverse relaxation times. Due to the nonlinearity of the Bloch equations, the superposition of the two radio frequency fields, $B_{\rm p}cos(\omega_{\rm p})+B_{\rm DM}cos(\omega_{\rm DM})$, produces an on-resonance ($\gamma B_{0}-\omega_{\rm DM}=0$) beat-note modulation of $m_{z}$,
@@
where $\omega_{\rm D}=|\omega_{\rm p}-\omega_{\rm DM}|$ being the detection frequency satisfying the condition $\omega_{\rm D}^{-1}<\min\left\{T_{1},T_{2}\right\}$. Eq. 8 holds at resonance and for dark matter frequency around $\rm 1~GHz$ $(\omega_{L}=\omega_{\rm DM}=\rm GHz)$. While a stronger pump field leads to larger amplification, the system must be kept far from saturation $B_{\rm p}\ll 1/\gamma\sqrt{T_{1}T_{2}}$. We choose $B_{\rm p}=0.1/\gamma\sqrt{T_{1}T_{2}}$ in our calculations. For further details on the detection scheme, see Appendix B.
@@
The integral is performed over the resonant part of the sample volume, where the electron spins are in magnetic resonance as defined by the condition $\omega_{\rm DM}=\omega_{L}\pm\delta\omega$ with linewidth given by $\delta\omega=\frac{1}{T_{2}}$ far from saturation. Near the zero-probe-field resonance regime, the entire sample can be within the resonance region [92], which gives a large force on the mechanical system, despite the weak field gradient. To determine the sensitivity of a MRFM setup to radio frequency field $B_{\rm DM}$, we now consider the contribution of various noise sources to the resonator motion.
Figure 2: Sensitivity plot of axion-electron coupling strength $g_{aee}$ (left axis) and effective magnetic field (right axis) as a function of axion mass $m_{\rm DM}$ (lower axis) and corresponding Compton frequency $f_{\rm DM}$ (top axis). Our results for MRFM based axion search are given in red (membrane) and blue (cantilever). For each resonator, a 100 day run at a single frequency is shown and a 100 day scan of minute long runs at neighboring frequencies.The QCD axion band is shown in yellow and existing constraints from astrophysical [28], space [49, 9] and laboratory experiments [94, 105] are shown in gray. Also shown are narrowband constraints from other magnetic resonance based searches in green, labeled QUAX [36, 35], UWA [44] and QND [56]. The top right panel shows the finer details of a single sensitivity line including the two sidebands at $\omega_{p}\pm\omega_{m}$. The axion sensitivity region can be moved easily by varying the pump field frequency $\omega_{p}$. The bottom right panel shows the contributions from the back action noise (purple), thermal noise (yellow), and imprecision noise (green) to the total noise equivalent force (red).
Figure 3: Sensitivity plot of dark photon kinetic coupling parameter $\epsilon$ as a function of dark photon mass $m_{\rm DM}$ (lower axis) and corresponding Compton frequency $f_{\rm DM}$ (upper axis). Our results for MRFM based axion search are given in red (membrane) and blue (cantilever). For each resonator, a 100 day run at a single frequency is shown and a 100 day scan of minute long runs at neighboring frequencies.The gray region represents constraints from astrophysical and cosmological observations assuming the dark photon makes up majority of dark matter, while dark green and light green are constraints from haloscope searches and rescaled limits from axion searches. See [29] for details of various constraints on the dark photon kinetic coupling.
@@
where the dark matter coherence time is $\tau_{c}\equiv\tau_{\rm DM}=\frac{c^{2}}{v_{\rm DM}^{2}f_{\rm DM}}=\frac{10^{6}}{f_{\rm DM}}$ for couplings proportional to the field and $\tau_{c}\equiv\tau_{\nabla a}=0.68\ \tau_{\rm DM}$ for couplings proportional to the gradient of the axion field $a$ [14]. Here, NEF denotes the noise equivalent force, which corresponds to a signal to noise ratio of 1.
@@
DM Parameter Space.– Before providing specific details of the device used for our estimates, we now overview existing constraints on the ULDM parameter space to work out the required magnetic field sensitivity. The parameter space for axion-electron coupling with existing laboratory and astrophysical constraints is shown in Fig. 2. The most stringent constraint on the axion-electron coupling at $g_{ae}\sim 10^{-13}~\rm eV$ stems from the lack of observed reduction in the tip of the red giant branch (TRGB) luminosity in globular clusters, where axion bremsstrahlung from electrons would otherwise cause significant energy loss and cooling [28]. Weaker constraints arise from solar axion searches, including limits on axion-induced solar energy losses from measurements of the neutrino flux [49] and null detection of electron recoils from solar neutrinos [9]. The best current laboratory constraint in the relevant mass range is due to torsion pendulum measurements of dipole-dipole forces between electrons [94] at $7\times 10^{-9}~\rm eV$, while measurement of the anomalous magnetic moment of electrons [105] also gives weaker constraints.
Three major milestones in the search for axions via their electron coupling are probing the coupling better than torsion pendulum limits [94], observations of tip of the red giant branch (TRGB) luminosity in globular clusters [28], and probing the QCD axion bands for DFSZ and KSVZ-like axions. These correspond to magnetic field strengths of $3\times 10^{-16}~\rm T$, $5\times 10^{-21}~\rm T$, and $2\times 10^{-24}~\rm T$, respectively (see Fig. 2). The best current narrowband experimental constraints are from magnetic axion haloscopes that rely on detecting the axion-magnon coupling [36, 35, 44, 56]. A number of proposals relying on optimistic advancements in technology have been proposed using QND measurements of axion-magnon conversion [34, 72, 17, 31, 32]. Our MRFM proposal for axion searches is similar in technical requirements and detection sensitivity to the best existing laboratory constraints, including QUAX [36, 35]. Unlike the aforementioned proposals, it does not require major technological advancements and the experiment can be performed with current technology.
@@
Using the parameters given in Table 1 and Eq. 12, we find a minimum resonant magnetic field sensitivity at $\rm GHz$ frequency of about $1.6\times 10^{-17}$ T ($8.4\times 10^{-19}$ T) for a minute (100 days) of integration time with a membrane and $2.0\times 10^{-17}$ T ($1.0\times 10^{-18}$ T) for a minute (100 days) of integration time with a cantilever. Using Equations 1 and 4, we obtain minimum values that can be probed by our setup of the axion-electron coupling and the dark photon kinetic mixing parameters, respectively. These are plotted in Figs. 2 and 3. For the axion-electron coupling, an MRFM experiment can search for axions beyond existing laboratory constraints. For the dark photon kinetic mixing, this system allows for filling in gaps between existing haloscope constraints.
@@
Length and Width $2.5~\rm mm$
Height $0.25~\rm mm$
Table 1: Experimental parameters used in our calculations of the sensitivity of MRFM to dark matter-induced magnetic fields. The values for the resonators are chosen in the range of parameters from [84, 93, 106] for membrane and [95] for cantilever.
The bottom right panel of Fig. 2 shows contributions to the noise equivalent force spectral density from back action noise (purple), thermal noise (yellow) and imprecision noise (green). The top right panel shows the corresponding plot for sensitivity to the axion-electron coupling, showing zoomed in part of both sidebands $\omega_{p}\pm\omega_{m}$. This two-sideband feature comes from the frequency up conversion scheme and has been seen by, for example, the AURIGA detector [25]. Our proposed experiment is limited by thermal noise over a relatively broad band with a linewidth of around $100~\rm Hz$ (see Fig. 5 in Appendix C).
@@
One way to improve sensitivity is to increase the force on the resonator by using a larger sample. In free space, the sample size is primarily limited by radiation damping via magnetic dipole emission. A resonant cavity can be used to avoid radiation damping altogether [36], allowing for the use of large magnetized samples.
The constraints presented here involve just a minute of integration time on existing technology, with the entire experimental setup shown in figure 1 taking $\sim$1 cm3 of volume. It is possible to fit multiple such setups inside a dilution refrigerator, improving the sensitivity by $\sqrt{N_{e}}$, for $N_{e}$ experimental setups, or scanning over a much larger bandwidth. A higher mechanical quality factor, as demonstrated on several platforms, can allow one to probe even weaker ULDM couplings. There is also potential for further multiplexing by using more magnetized samples, which only requires one to adjust the applied bias field to widen the resonance region. Furthermore, for axion-electron couplings in particular, flux concentrators can be used to amplify the signal [46]. Finally, our proposal involves standard quantum limit readout, a limit that has already been circumvented in several experimental platforms, e.g. [102, 60, 71].
Conclusion and outlook. – While most MRFM research has pushed towards detecting single spin flips, our work motivates exploration of the opposite regime, involving larger samples. Our approach involves a frequency scanning technique that can be adapted to other optomechanical searches for dark matter [70, 69, 55]. A similar MRFM based search can be conducted for axion-nucleon couplings using a nuclear spin sample, although the detection scheme may differ due to slower nuclear timescales. Our proposal for a magnetic resonance force microscope search for dark matter induced magnetic fields is just one thrust within the larger mechanical systems to search for fundamental physics theme [30, 73, 62, 4, 24, 8, 80, 6, 7]. Beyond dark matter, other interesting signals from gravitational waves, neutrinos, etc. can potentially be probed using optomechanical systems, paving the way for pushing forward technology and potentially discovering new physics.